OpenStudy (usukidoll):

(y+1)dx+(y-1)(1+x^2)dy = 0

OpenStudy (usukidoll):

the most frustrating separate variable one in my honest opinion. I did this like 10 times and got stuck

OpenStudy (anonymous):

Is this the original equation?

OpenStudy (usukidoll):

yeah. I'm gonna get an attachment to this

OpenStudy (anonymous):

I'd think this is the original equation:\[ (y-1)(1+x^2)y'=-(y+1) \]

OpenStudy (anonymous):

Separating variables\[ -\frac{y-1}{y+1}y'=\frac{1}{1+x^2} \]

OpenStudy (anonymous):

Integrating both sides with respect to \(x\): \[ \int -\frac{y-1}{y+1}y'dx=\int \frac{1}{1+x^2}dx \]Remember that \(dy=y'dx\)

OpenStudy (usukidoll):

it's the y that's driving me nutssss...isn't an integral from the tables of integrals or something?

OpenStudy (anonymous):

\[ u=y+1,\quad y-1=u-2, \quad du=dy \]

OpenStudy (usukidoll):

where did that come from?

OpenStudy (usukidoll):

simple subsitution?

OpenStudy (anonymous):

You want to get rid of an ugly denominator.

OpenStudy (anonymous):

My thought process: any linear substitution is cheap and easy.

OpenStudy (usukidoll):

then I get a 1/u if u = y+1

OpenStudy (anonymous):

\[ \int\frac{u-2}{u}\;du \]

OpenStudy (usukidoll):